Perform the indicated operations. Write the answer in standard form, a+bi. (V2 + 10 i)(12 - 101) 0 0/0 х 5
Answer #1

Solution:

\\\left(\sqrt{2}+\sqrt{10}i\right)\left(\sqrt{2}-\sqrt{10}i\right)\\\\\mathrm{Apply\:complex\:arithmetic\:rule}:\quad \left(a+bi\right)\left(a-bi\right)=a^2+b^2\\\\a=\sqrt{2},\:b=\sqrt{10}\\\\=\left(\sqrt{2}\right)^2+\left(\sqrt{10}\right)^2\\\\=2+10\\\\\left(\sqrt{2}+\sqrt{10}i\right)\left(\sqrt{2}-\sqrt{10}i\right)=12